# FOCUS on Contamination: A Geospatial Deep Learning Framework with a Noise-Aware Loss for Surface Water PFAS Prediction

### v3

@@ Introduction
Based on discussions with domain experts and experimental validation (see supplement), we set the weights as $\alpha_{1}=0.4$, $\alpha_{2}=0.2$, $\alpha_{3}=0.1$, and $\alpha_{4}=0.3$. Pixels at ground truth sample points labeled 1 are assigned a probability of 1, reflecting complete confidence. However, for sample points labeled 0, we account for additional uncertainty due to certain PFAS compounds having method detection limits (MDLs) that exceed advisory thresholds [^1] In such cases, a compound may go undetected despite being present at a high-risk level. To address this, the corresponding noise mask value is down-weighted in proportion to the ratio of the advisory threshold to the MDL.
Model
Figure 5: Model overview: high-level representation of FOCUS
Our framework, FOCUS, is inspired by the Prithvi architecture (Blumenfeld 2023) and employs a masked autoencoder (MAE) approach for pretraining (Fig. 5). Instead of relying on Prithvi’s pretrained weights, we pretrain our model on derived geospatial data products (e.g., land cover and distance rasters), which we hypothesize offer more contextual information for capturing the environmental nuances critical to PFAS contamination prediction (see section Results). This pretraining uses masked autoencoding with a mean squared error loss. The pretrained encoder is then fine-tuned with a decoder head dedicated to PFAS prediction.
Loss Function
The loss function used in FOCUS combines the focal loss with noise mask weighting as follows:
$\mathcal{L}_{\text{CE}}(p_{i},y_{i})$ $=-\left[y_{i}\cdot\log(p_{i})+(1-y_{i})\cdot\log(1-p_{i})\right],$ (2)
$\mathcal{L}_{\text{focal}}(p_{i},y_{i})$ $=(1-p_{i})^{\gamma}\cdot\mathcal{L}_{\text{CE}}(p_{i},y_{i}),$ (3)
$\mathcal{L}_{\text{{FOCUS}}}$ $=\frac{1}{N}\sum_{i=1}^{N}\mathcal{L}_{\text{focal}}(p_{i},y_{i})\cdot M_{i}.$ (4)
@@
where $N$ is the total number of pixels, $y_{i}\in\{0,1\}$ is the ground truth label for pixel $i$, $p_{i}$ is the predicted probability of the positive class, and $M_{i}\in[0,1]$ is the corresponding noise mask value indicating confidence in the label.
The parameter $\gamma=2$ is used in the focal loss to focus on hard-to-classify examples, as this value showed best performance in previous studies introducing focal loss (Lin et al. 2018). Notably, class weights to address dataset imbalance are directly incorporated into the binary cross-entropy loss ($\mathcal{L}_{\text{CE}}$), removing the need for an explicit alpha term in focal loss. By combining focal loss and noise mask weighting, the loss function addresses label noise and class imbalance while emphasizing high-confidence pixels during training.
Experiments
Baselines
@@
  • Pollutant transport simulation: A scientific, process-based approach that models contaminant movement using hydrological principles (SWAT n.d.).
  • Landsat-based method: Uses the original Prithvi weights to predict PFAS from raw satellite imagery (Landsat 7 multispectral (U.S. Geological Survey n.d.)).
  • FESTA loss approach: An alternative state-of-the-art technique designed for sparse segmentation using sparse point data. (Hua et al. 2022).
  • Kriging: A geostatistical interpolation technique (Esri n.d.) that predicts contamination based on spatial relationships between points.
  • Random forest: Following (DeLuca et al. 2023), this model aggregates environmental features within a 5 km buffer around each sample (e.g., calculating the percentage of various land cover types) to predict contamination at individual points.
  • Rule-based approach: Assigns contamination labels based on predefined environmental heuristics, replicating our noise mask approach to compute a contamination probability (but without ($p_{\text{sample\_dist}}$)), which is then thresholded to yield binary predictions.
Training Configuration
We now describe our experimental setup, with more details in the supplement. We conducted experiments per year, allowing us to evaluate the model’s ability to fill spatial data gaps rather than forecast across time. We split each year’s data into 80% for training and 20% for testing, ensuring that training and testing occur within the same year’s distribution. To ensure geographic diversity and avoid spatial leakage, we further ensured that the test set was geographically disjoint from the training set.
A batch size of 4 and a learning rate of $5\times 10^{-4}$ were adopted to preserve the stability of our pretrained backbone, minimizing the risk of large, destabilizing updates. In addition, we employed AdamW (Loshchilov and Hutter 2019) as the optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.999$. We tuned the learning rate and batch size empirically by monitoring training stability. The final values were selected based on consistent convergence and favorable test-time metrics across years. Furthermore, we used a custom learning rate scheduler with an initial warmup phase followed by polynomial decay, which helped stabilize convergence and ensured that the model gradually adapted to the limited data.
Figure 6: Comparison of performance across different methods and years. Results are averaged over 3 random seeds for FOCUS and applicable baselines (Landsat-based, FESTA loss, rule-based).
Ablation Study
@@
  • Noise Weights: We evaluate the impact of incorporating noise masks into the FOCUS loss function. With noise masks, pixel-wise confidence scores scale focal loss to emphasize high-confidence pixels; without them, the loss reduces to standard focal loss (Lin et al. 2018).
@@
While the primary focus of this work is on PFAS contamination, we also explored the generalizability of our framework to land cover classification, and robustness through a consistency analysis detailed in the supplement.
Metrics
@@
The reported metrics include accuracy, Intersection over Union (IoU), F-score, precision, and recall. Performance metrics in this study are reported exclusively from evaluations using only ground truth samples. In addition, we assess timing efficiency as a metric to evaluate the model’s computational scalability.
Results
FOCUS Outperforms Existing Baselines
Main Results
We first compared FOCUS against six alternative approaches for PFAS contamination prediction as shown in Fig. 6. Overall, FOCUS achieves better performance across all the years and most metrics, aside from accuracy in 2019 and 2022, and precision in 2022. The Landsat-based method yielded moderate results, suggesting that geospatial data products are more effective for capturing PFAS contamination patterns than relying solely on raw satellite data. The random forest approach, despite achieving acceptable accuracy, suffered from lower F-score, precision, and recall, likely due to the limited spatial context from per-point feature aggregation. Similarly, although Kriging provided decent accuracy and precision, it struggled with recall, indicating difficulties in detecting all instances of contamination, and pollutant transport simulation demonstrated weak overall performance. The FESTA loss approach showed suboptimal performance in our experiments. This may be attributed to our dataset’s exclusive use of point data and the complex dynamics of PFAS contamination, which may challenge the spatial continuity assumptions of the FESTA loss. While rule-based masks, designed to approximate contamination using domain-specific heuristics only, can provide a reasonable estimate (and prove useful for computing noise masks, see section Methodology), they fall short in accuracy compared to the data-driven approach. This finding underscores the importance of incorporating domain expertise and AI for more reliable PFAS prediction.
Time Efficiency
We followed the methodology of DeLuca et al. (DeLuca et al. 2023), which uses similar environmental features, dataset, and a random forest model, as a baseline for efficiency comparisons. Table 2 compares feature extraction and inference times between the ML method and our FOCUS approach over a 1 km2 area and in Northern Michigan (NM) (approximately 44,000 km2).
@@ Results
For feature extraction, the ML baseline required aggregating environmental features within a 1 km buffer per point, taking an average of 1.2 minutes per point and nearly 2 days over the NM region. In contrast, FOCUS processes full 1 km2 raster patches in 4.7 seconds on average, reducing extraction time to 3.2 hours for NM. Inference times were comparable: 7.5 seconds for ML vs. 13 seconds for FOCUS over NM. Table 2 summarizes these results. All times were averaged over 10 runs. Feature extraction was performed on an Intel Xeon Gold 6154 CPU (3.0 GHz), and inference used an AMD Ryzen Threadripper PRO 7985WX with an NVIDIA RTX 6000 Ada GPU.
@@
Table 2: Comparison of performance for feature extraction and inference using random forest and FOCUS over 1 km2 area and Northern Michigan (NM).
FOCUS Applies in the Real World
We next conducted two complementary validation efforts of predicted contamination maps of new regions – beyond regions surrounding ground truth sample points in the train and test sets – using field-collected data (see Table 1). First, our research team collected surface water samples from multiple local sites in Michigan. To reiterate, these represent previously unsampled regions and provide an independent means of assessing model generalization beyond our test set. Notably, as shown in Fig. 7, the model correctly predicted 8 collected points in this evaluation. Technical details on sample analysis are available in the supplement.
Figure 7: Real-world validation using surface water samples collected by our team. Newly sampled locations (green) are overlaid on model predictions, with red indicating areas of high predicted contamination, and dark green indicating areas of lower predicted contamination.
Second, we evaluated our 2022 fish tissue model on the publicly available MPART fish dataset from 2022 (Michigan Department of Environment, Great Lakes, and Energy 2025), which includes PFAS measurements in fish tissue from across Michigan, collected independently from our fish train and test sets. Results are comparable to Fig. 6.
@@
MPART Fish Samples 84 $\pm$ 2 % 61 $\pm$ 1 % 74 $\pm$ 1 % 74 $\pm$ 1 % 73 $\pm$ 2 %
Table 3: FOCUS performance on fish tissue samples collected independently of training data. Results averaged over 3 random seeds.
FOCUS Noise Masks Improve Performance
@@
Accuracy 79 $\pm$ 2 88 $\pm$ 1 89 $\pm$ 1 89 $\pm$ 1 84 $\pm$ 3 86 $\pm$ 2
IoU 56 $\pm$ 3 65 $\pm$ 2 67 $\pm$ 2 71 $\pm$ 2 65 $\pm$ 3 68 $\pm$ 2
F-score 70 $\pm$ 3 77 $\pm$ 2 79 $\pm$ 1 82 $\pm$ 2 77 $\pm$ 3 80 $\pm$ 2
Precision 68 $\pm$ 2 83 $\pm$ 2 78 $\pm$ 2 78 $\pm$ 2 74 $\pm$ 3 76 $\pm$ 2
Recall 74 $\pm$ 2 73 $\pm$ 2 80 $\pm$ 1 89 $\pm$ 1 91 $\pm$ 1 92 $\pm$ 2
Table 4: Ablation showing impact of noise-aware loss (FOCUS) vs. standard focal loss. Results averaged over 3 random seeds. Bold values denote the best result per year.
Table 4 evaluates the effect of incorporating noise masks in the FOCUS loss computation, compared to focal loss only (i.e., removing $M_{i}$ from Equation 4). We find that the addition of noise masks consistently improves performance, though FOCUS may sometimes favor precision over sensitivity. This is due to emphasizing high-confidence pixels and down-weighting uncertain ones during training.
Note that this ablation is conducted at the $256\times 256$ resolution, which was selected based on the ablation study comparing it with $512\times 512$ patches, detailed in the supplement. The smaller resolution consistently performed better, likely due to reduced label noise propagation across larger contexts. We did not evaluate patch sizes smaller than $256\times 256$, as excessively small patches can compromise accuracy and increase variability (Lechner et al. 2009).
FOCUS has Robust Results

### v4

@@ 3.4. Physically-Informed Pixel Confidence $M_{i}$ Details
We perform an ablation study on the 2008 PFAS dataset using a grid search over 24 noise-mask configurations derived from four inputs: dischargers, land cover, flow direction, and distance to sample points. We vary their relative weights and evaluate segmentation performance across multiple metrics. The best configuration assigns 40% to dischargers, 20% to land cover, 10% to sample distance, and 30% to flow direction. These results align with domain expert guidance that emphasizes the importance of industrial activity in driving contamination patterns. Pixels at ground truth sample points labeled 1 are assigned a probability of 1, reflecting complete confidence. However, for sample points labeled 0, we account for additional uncertainty due to certain PFAS compounds having method detection limits (MDLs) that exceed advisory thresholds [^2] In such cases, a compound may go undetected despite being present at a high-risk level. To address this, the corresponding noise mask value is down-weighted in proportion to the ratio of the advisory threshold to the MDL. Full performance results for all configurations and additional details on the environmental priors are reported in the appendix.
3.5. FOCUS Loss
We scale focal loss by $M_{i}$ to emphasize trustworthy pixels and hard cases:
(4) $\mathcal{L}_{\text{FOCUS}}=\frac{1}{N}\sum_{i=1}^{N}M_{i}(1-p_{i})^{\gamma}\Big[-y_{i}\log p_{i}-(1-y_{i})\log(1-p_{i})\Big]$
@@ 3.5. FOCUS Loss
where $N$ is the total number of pixels, and $p_{i}$ is the predicted probability of the positive class. Key components:
  • $M_{i}$ down-weights uncertain pixels (noise-robust)
  • $(1-p_{i})^{\gamma}$ up-weights hard examples (focal modulation)
Proposition 1 (Noise-Aware Surrogate Objective) Let $z_{i}\in\{0,1\}$ denote the latent contamination state and $y_{i}\in\{0,1\}$ the observed expanded label under a pixel-specific asymmetric flip noise rate $\eta_{i}<\tfrac{1}{2}$, such that $\Pr(y_{i}=1\mid x_{i})=\eta_{i}+(1-2\eta_{i})\,\pi_{\theta}(x_{i}).$ Then there exist coefficients $a_{i},b_{i}$ such that the noisy negative log-likelihood satisfies $\ell_{\text{noisy}}(p_{i},y_{i})\leq a_{i}\cdot\mathrm{CE}(p_{i},y_{i})+b_{i}.$ where $\mathrm{CE}$ denotes the standard cross-entropy loss. Setting $a_{i}=M_{i}$, where $M_{i}$ increases as $\eta_{i}$ decreases, yields a valid surrogate objective whose expectation upper-bounds the $\ell_{\text{noisy}}(p_{i},y_{i})$ and serves as a lower-bound surrogate to the clean log-likelihood.
Proof sketch. Under the asymmetric flip noise model with pixel-specific noise rate $\eta_{i}<\tfrac{1}{2}$, the noisy likelihood satisfies
$\Pr(y_{i}=1\mid x_{i})=\eta_{i}+(1-2\eta_{i})\,p_{i},$
where $p_{i}=\pi_{\theta}(x_{i})$. The corresponding negative log-likelihood is
$\ell_{\text{noisy}}(p_{i},y_{i})=-\log\!\left(\eta_{i}+(1-2\eta_{i})\,p_{i}\right).$
Since the logarithm is concave, for any fixed anchor $p_{0}\in(0,1)$ there exist coefficients $a_{i},b_{i}$ such that
$\ell_{\text{noisy}}(p_{i},y_{i})\leq a_{i}\cdot\mathrm{CE}(p_{i},y_{i})+b_{i},$
We then apply the standard focal modulation $(1-p_{i})^{\gamma}$, where $\gamma\geq 0$ controls the focus on hard examples, to the CE term.
If we set $a_{i}=M_{i}$, where $M_{i}$ increases as $\eta_{i}$ decreases (i.e., higher label trust), then $\mathcal{L}_{\text{FOCUS}}$ is a valid upper-bounding surrogate of $\mathbb{E}[\ell_{\text{noisy}}]$ (up to additive constants).
Consequently, minimizing $\mathcal{L}_{\text{FOCUS}}$ maximizes a lower bound on the (noisy) log-likelihood, and—under the alignment assumption on $M_{i}$—serves as a principled surrogate for learning from the underlying clean labels while encoding physically grounded priors rather than learned heuristics. Further details and a formal derivation are provided in the appendix.
3.6. Model
Our framework, FOCUS, is inspired by the Prithvi architecture (Blumenfeld, 2023) and employs a masked autoencoder (MAE) approach for pretraining (Fig. 4). Instead of relying on Prithvi’s pretrained weights, we pretrain our model on derived geospatial data products (e.g., land cover and distance rasters), which we hypothesize offer more contextual information for capturing the environmental nuances critical to PFAS contamination prediction (see Section 5).
4. Experiments
4.1. Baselines
@@
  • Pollutant transport simulation: A scientific, process-based approach that models contaminant movement using hydrological principles (SWAT, nd). Further details on implementation in appendix.
  • Landsat-based method: Uses the original Prithvi weights to predict PFAS from raw satellite imagery (Landsat 7 multispectral (U.S. Geological Survey, nd)).
  • FESTA loss approach: A state-of-the-art method for sparse segmentation from point supervision. (Hua et al., 2022).
  • Kriging: A geostatistical interpolation method (Esri, nd) that predicts contamination based on spatial relationships between points. Further details on implementation in appendix.
  • Random forest: Closely following the methodology of (DeLuca et al., 2023), this model aggregates environmental features within a 5 km buffer around each sample (e.g., calculating the percentage of various land cover types) to predict contamination at individual points.
  • Rule-based approach: Assigns contamination labels based on predefined environmental heuristics only, replicating our noise mask approach to compute a contamination probability (but without ($p_{\text{sample\_dist}}$)), which is then thresholded to yield binary labels.
4.2. Training Configuration
We now describe our experimental setup, with additional details in the appendix. We conduct experiments per year to evaluate the model’s ability to fill spatial data gaps, an important stakeholder need. We do not forecast across time as longitudinal ground truth is currently too sparse and inconsistent for reliable temporal supervision. For each year, we split the data into 80% training and 20% testing, with geographically disjoint regions to avoid spatial leakage and ensure diversity.
We use a batch size of 4 to preserve the stability of the pretrained backbone and optimize with AdamW (Loshchilov and Hutter, 2019) ( $\beta_{1}=0.9$ and $\beta_{2}=0.999$). A custom learning-rate schedule with warmup followed by polynomial decay stabilizes convergence and enables gradual adaptation to limited data. Hyperparameters were selected based on consistent convergence.
Figure 6. Comparison of performance across different methods and years. Results are averaged over 3 random seeds for FOCUS and applicable baselines (Landsat-based, FESTA loss, rule-based). IoU is not applicable to methods that do not produce dense pixel-level segmentations (e.g., Kriging and Random Forest), and is therefore omitted for those approaches.
4.3. Ablation Study
@@
  • Noise Weights: We evaluate the impact of incorporating noise masks into the FOCUS loss function. With noise masks, pixel-wise confidence scores scale focal loss to emphasize high-confidence pixels; without them, the loss reduces to standard focal loss (Lin et al., 2018).
@@
While the primary focus of this work is on PFAS contamination, we also explored the generalizability of our framework to land cover classification, and robustness through a consistency analysis detailed in the appendix.
4.4. Metrics
@@
The reported metrics include accuracy, Intersection over Union (IoU), F-score, precision, and recall. Performance metrics in this study are reported exclusively from evaluations using only ground truth samples. In addition, we assess timing efficiency (in seconds) as a metric to evaluate the model’s computational scalability.
For threshold-dependent metrics, we select the decision threshold using out-of-fold predictions on the training data. Specifically, we perform $K$-fold cross-validation on the training set, sweep thresholds to maximize F-score on the resulting out-of-fold predictions, and apply the selected threshold to the held-out test set. We use $K=5$ folds and sweep 200 evenly spaced thresholds in $[0,1]$. Thresholds are selected independently for each method.
5. Results
5.1. FOCUS Outperforms Existing Baselines
5.1.1. Main Results
We first compared FOCUS against six alternative approaches for PFAS contamination prediction, as shown in Fig. 6. Overall, FOCUS provides the most consistent performance across years and metrics.. While several baselines achieve competitive accuracy, according to domain experts, accuracy alone is insufficient in this highly imbalanced setting, making precision–recall trade-offs more informative. In particular, Kriging and the FESTA loss achieve comparable accuracy in some years and show reasonable performance across metrics, highlighting the strength of spatial interpolation and structure-aware learning in this setting. The Landsat-based model yields moderate results, suggesting that satellite features alone are insufficient without explicit geospatial context. The random forest approach, despite achieving acceptable accuracy, suffered from lower F-score, precision, and recall, likely due to the limited spatial context from per-point feature aggregation. Pollutant transport simulation and the rule-based approach provide useful physics- and heuristic-driven references (the latter also serving as a component in our noise modeling), but they lag behind data-driven methods in overall predictive balance. Taken together, these results indicate that FOCUS consistently achieves a favorable trade-off between precision and recall across years, which is critical for robust large-scale environmental screening.
5.1.2. Time Efficiency
@@
We compare FOCUS against a random forest baseline (following DeLuca et al. (DeLuca et al., 2023)) using similar environmental features and datasets. Table 2 reports feature extraction and inference times over a 1 km2 area and all of Northern Michigan (NM, $\sim$44,000 km2). The ML baseline requires aggregating features within a spatial buffer per point, resulting in significantly higher extraction costs at scale. In contrast, FOCUS processes full raster patches efficiently, reducing feature extraction time from days to hours over NM, while inference times remain comparable between methods. Feature extraction was performed on an Intel Xeon Gold 6154 CPU (3.0 GHz), and inference used an AMD Ryzen Threadripper PRO 7985WX with an NVIDIA RTX 6000 Ada GPU.
Table 2. Comparison of performance for feature extraction and inference using random forest and FOCUS over 1 km2 area and Northern Michigan (NM). Results averaged over 10 runs.
@@
5.2. FOCUS Applies in the Real World
We next conducted two complementary validation efforts of predicted contamination maps of new regions – beyond regions surrounding ground truth sample points in the train and test sets – using field-collected data (see Table 1). First, our research team collected surface water samples from multiple local sites in Ann Arbor, Michigan, in mid-2025. These sites represent previously unsampled regions and provide an independent assessment of model generalization beyond our test set under sparse and biased sampling, as testing is typically limited due to cost and feasibility constraints. Notably, as shown in Fig. 7, the model, trained on 2024 surface water data from MPART, correctly identified the 8 collected points as highly contaminated, demonstrating FOCUS’s ability to generalize under severe label scarcity. We are actively pursuing larger-scale validation with stakeholders and regulators. Technical details on sample analysis are provided in the appendix.
Figure 7. Real-world validation using surface water samples collected by our team. Newly sampled locations (green) are overlaid on model predictions, with red indicating areas of high predicted contamination, and dark green indicating areas of lower predicted contamination.
Second, we evaluated our 2019 fish tissue model in Table 3 on the publicly available MPART fish dataset from 2019 (Michigan Department of Environment, Great Lakes, and Energy, 2025), which includes PFAS measurements in fish tissue from across Michigan, collected independently from our fish train and test sets. Results are comparable to Fig. 6.
Qualitatively, Fig. 8 shows grayscale probability maps with red overlays indicating confident positive predictions in low-trust regions; the focal baseline produces many such responses, whereas FOCUS largely suppresses them, yielding more conservative behavior under unreliable supervision.
(a) Pixelwise confidence $M_{i}$ (white = higher trust)
(b) Focal probability map
(c) FOCUS probability map
Figure 8. Grayscale PFAS probabilities (white = higher). Red overlays indicate overconfident positives in low-trust regions $(p_{i}\geq 0.50,\;M_{i}\leq 0.30)$. FOCUS suppresses such responses, while the focal baseline produces many.
Table 3. FOCUS performance on fish tissue samples collected independently of training data. Results averaged over 3 random seeds.
@@ 5.2. FOCUS Applies in the Real World
MPART Fish Samples 85 $\pm$ 2 % 60 $\pm$ 1 % 72 $\pm$ 1 % 75 $\pm$ 1 % 70 $\pm$ 2 %
5.3. FOCUS Noise Masks Improve Performance
Table 4. Ablation showing impact of noise-aware loss (FOCUS) vs. standard focal loss. Results averaged over 3 random seeds. Bold values denote the best result per year.
@@ 5.3. FOCUS Noise Masks Improve Performance
Accuracy 37 $\pm$ 2 73 $\pm$ 1 62 $\pm$ 1 82 $\pm$ 1 55 $\pm$ 3 88 $\pm$ 2
IoU 22 $\pm$ 3 48 $\pm$ 2 41 $\pm$ 2 60 $\pm$ 2 36 $\pm$ 3 67 $\pm$ 2
F-score 36 $\pm$ 3 61 $\pm$ 2 57 $\pm$ 1 73 $\pm$ 2 53 $\pm$ 3 79 $\pm$ 2
Precision 54 $\pm$ 2 61 $\pm$ 2 63 $\pm$ 2 70 $\pm$ 2 63 $\pm$ 3 77 $\pm$ 2
Recall 55 $\pm$ 2 64 $\pm$ 2 78 $\pm$ 1 85 $\pm$ 1 74 $\pm$ 1 81 $\pm$ 2
Table 4 evaluates the effect of incorporating noise masks in FOCUS, compared to focal loss alone (i.e., removing $M_{i}$ from Eq. 4). Adding noise masks consistently improves performance by emphasizing high-confidence pixels and down-weighting uncertain ones during training. This ablation uses a $256\times 256$ resolution, selected based on the patch-size study in the appendix, where it outperformed $512\times 512$, likely due to reduced label-noise propagation over large contexts. We did not evaluate smaller patches, as overly small contexts can degrade accuracy and increase variability (Lechner et al., 2009).
5.4. FOCUS has Robust Results
To evaluate generalization to unseen regions, we perform spatial cross-validation by holding out sets of U.S. states (Table 5). Specifically, three disjoint groups of five states each are used as test sets, with the remaining states used for training, ensuring no spatial overlap and mimicking deployment in unsampled areas. FOCUS consistently outperforms focal loss across all years, demonstrating stronger spatial generalization and robustness under distribution shifts.
We also assess spatial robustness via a consistency analysis across all 49 continental U.S. states by comparing predictions on overlapping patches at two overlap scales ($\times56$ and $\times156$ pixels). We observe high agreement (above 93% on average), indicating locally consistent predictions; full details are provided in the appendix. Results on model calibration (i.e., Expected Calibration Error (ECE) (Guo et al., 2017)) are also provided in the appendix, showing low ECE values ($\leq 0.1$) and indicating that predicted probabilities are well-aligned with observed outcomes.
Table 5. Cross-state generalization results ($\text{mean}\pm\text{std},$ over the disjoint state splits). Bold values denote the best result per year.